Good Day,
I have a question: The present age of the father is 8 times his son's age. After 2 years, his age will be 6 times his son's age. What are their present ages?

Answer by richwmiller(17219)

f=8s
f+2=6*(s+2)
f+2=6s+12
f=6s+10
8s=6s+10
2s=10
s=5
f=40
check
42=6*(7)
42=42
ok